OpenStudy (anonymous):

what is a solution of x^2+6x = -5 a. x = -6 b. x = -1 c. x = 1 d. x = 6

OpenStudy (welshfella):

x^2 +6x + 5 = 0

OpenStudy (anonymous):

x^2+6x+5=0 (x+1)(x+5)=0

OpenStudy (welshfella):

you need 2 numbers whose product is 5 and sum = 6

OpenStudy (welshfella):

@amirreza1870 - plz just dont give the answer - read the Code of Conduct

OpenStudy (welshfella):

It doesn't help the questioner learn how to solve problems

OpenStudy (anonymous):

is it 6? @amirreza1870

OpenStudy (amorfide):

re arrange the expression in the form ax^2 + bx + c = 0 so rearrange to get x^2 +6x + 5 = 0 now you want to find two numbers which multiply to give 5 but add up to give 6 so write out your multiples of 5 then you will write what you think in the form (x+a)(x+b) you can check if you are correct by multiplying out

OpenStudy (anonymous):

its 1

OpenStudy (anonymous):

@amorfide

OpenStudy (amorfide):

nope

OpenStudy (amorfide):

x+1=0 is a solution but x is not equal to 1

OpenStudy (anonymous):

-1? @amorfide

OpenStudy (amorfide):

yes